OpenStudy (anonymous):

Solve for x: √x+√(x-20)=10 NEED HELP!!!!

OpenStudy (anonymous):

Ok square both sides: \[2x-20+2\sqrt(x(x-20))=100\] Isolate the radical: \[2\sqrt{x(x-20)} = 120 -2x\] Square again: \[4(x-20)x=(120-2x)^2\] Expand terms: \[4x^2-80x=4x^2-480x+14400\] Subtract to simplify: \[400x-14400=0\] Factor out and divide x-36=0 So the answer is 36.

OpenStudy (anonymous):

Thank You! You're good in math.. I have more questions.. Can you help me to solve it?? (:

OpenStudy (anonymous):

If I can have medals, ya

OpenStudy (anonymous):

have you received mine?